OpenStudy (shamil98):

Let log P/N = 5 , and log M/N = 9 , What is the relationship between P and M?

OpenStudy (ash2326):

Subtract log M/ N by log P/N

OpenStudy (anonymous):

is log the natural log?

OpenStudy (shamil98):

no the log as in 10 .

OpenStudy (shamil98):

How do I do that ash? My knowledge with logs is little ..

OpenStudy (ash2326):

ok \[\log \frac M N - \log \frac P N =9-5\] Can you simplify this a bit? maybe the right side

OpenStudy (shamil98):

Yeah, 4 on the right side

OpenStudy (ash2326):

According to subtraction property of log \[\log A -\log B= \log \frac A B\] Can you use that here? @shamil98

OpenStudy (shamil98):

log M/N = log M - log N log P/N = log P - log N

OpenStudy (ash2326):

treat M/ N as A P/N as B then use it to reduce two terms to one

OpenStudy (ash2326):

never mind that would work as well, substitute your expansion and see if any terms get cancelled.

OpenStudy (anonymous):

\[ \frac P N = 10^5 \\ \frac M N = 10^9 \\ \frac {\frac P N} {\frac M N}= \frac {10^5}{10^9}\\ \frac {P}{M}=\frac{1}{10^4} \]

OpenStudy (shamil98):

Log M - log N - ( log P - log N ) = 4 Log M - log N - log P + log N = 4 Log M - Log P = 4

OpenStudy (ash2326):

yes, now use the sub property to combine log m and log p

OpenStudy (shamil98):

log M/P = 4

OpenStudy (ash2326):

@eliassaab Sir please don't directly handover the answers to the asker. It won't help them.

OpenStudy (ash2326):

good shamil. Do you know how log is defined?

OpenStudy (shamil98):

Nope.. could you explain?..

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Logarithm

OpenStudy (ash2326):

ok, so we have \[\log_A B=C\] this implies \[B=A^C\] A=base that's how log is defined Suppose we have \[\log_2 4=C\] so \[2^C=4\] or \[C=2\] Do you follow?

OpenStudy (shamil98):

Yeah.

OpenStudy (ash2326):

can you apply it here?

OpenStudy (shamil98):

log M/P = 4 \[\log_{10} \frac{ M }{ P } = 4\] \[\large \frac{ M }{ P } = 10^4\] ?

OpenStudy (ash2326):

yes, good work. You can simplify it

OpenStudy (shamil98):

\[\large \frac{ M }{ P } = 10000\] \[\large M = 10000P\]

OpenStudy (ash2326):

yes, that's right

OpenStudy (shamil98):

thanks, for the explanation. I appreciate @eliassaab your effort, I just needed an explanation on how to get to there.